49x^2-25=0
49x^2-25=0
(7X+5)(7X-5) =0
7X+5=0 OR 7X-5=0
X=-5/7 ,X=5/7

Found 3 solutions by solver91311, mananth, chiexpert:
Answer by solver91311(24713)



Did you bother to read the instructions for posting on this site? I specifically refer to the part where it says: "Please try to ask your question clearly and fully so that it MAKES SENSE. We are not mind readers. We do not have your textbook in front of us. All we know is what you wrote." Or did you simply think that such mundane nonsense couldn't possibly apply to YOU?

Read these RULES before submitting your question!
Here's your opportunity to send a question to our free, volunteer tutors. If your question makes sense, it will be answered by one of Algebra.Com's tutors.

An email will be sent to you if you leave your email address (it is confidential), when the problem is solved.

Have you seen these solvers: linear equation and system, quadratic equation, Univeral Simplifier and equation solver? They (and dozens of other solvers) can save your time and give you an instant answer with very detailed explanation.
Daily limit of 5 problems per day. All tutors on algebra.com are FREE.

One problem per submission. Do not dump your whole homework in one question.
If you can, show what you have tried. Tutors like students who try. If a tutor gives you credit for work shown, that problem would not count towards your daily limit of 5 per day.

Please try to ask your question clearly and fully so that it MAKES SENSE. We are not mind readers. We do not have your textbook in front of us. All we know is what you wrote.

Do not use slang or strange abbreviations. Spell every word as well as you can. This improves your chances that your question will be answered.

When your problem is solved and you view the solution, use the thank you form to say thank you to the tutor. it is very important!

If you like someone's answers, you can speak to them about getting their private tutoring.
Example of a BAD QUESTION: ``Ineed ur help on binomils, plz plz I am dying!''

Example of a GOOD QUESTION: `' What is the slope of a binomial 2x+3''

John

My calculator said it, I believe it, that settles it


Answer by mananth(16946)

49x^2-25=0
(7x)^2-(5)^2=0
(7x+5)(7x-5)=0
x=-5/7 ,OR x=5/7

Answer by chiexpert(48)

49x^2-25=0
(7X+5)(7X-5) =0
7X+5=0 OR 7X-5=0
X=-5/7 ,X=5/7